Find and label the vertex and the line of symmetry.
f(x)=(1/5)x^2
Vertex= _______
Line of Symmetry = ________
THANKS

Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1862)

f(x)=(1/5)x^2 is in form y=a(x-h)^2+k ---> (h,k) vertex and x=h line of symmetry

h=0, k=0

Vertex: (0,0)
Line of Symmetry: x = 0

Answer by stanbon(75887)

Find and label the vertex and the line of symmetry.
f(x)=(1/5)x^2
Vertex= _______
Line of Symmetry = ________
---------------------
Vertex form:
(y-0) = (1/5)(x-0)^2
Vertex: (0,0)
--------------------------
Axis of Symmetry: x=0
===========================
Cheers,
Stan H.